# promptbox[1][] colframe=black, colback=white, sharp corners, boxrule=1pt, boxsep=8pt, left=0pt, right=0pt, top=0pt, bottom=0pt, #1

### v1

@@ 6.6.3. Experiment design
The goal of the experiment design presented in this paper was to validate the feasibility of our approach in stimulating changes in system outcomes and invoking user empathy. As a result, the experiment did not aim to directly measure the extent to which users acquired privacy knowledge or the subsequent shifts in their behavior as a consequence of our approach. Although the qualitative findings of the interviews with the participants suggest a strong potential of our approach to achieve these two goals, follow-up experiments are needed to validate such hypotheses. In the future, we plan to conduct pre-tests and post-tests with users between their usage of our system to assess their acquisition of privacy knowledge, as well as longitudinal deployment studies for measuring changes in users’ privacy behaviors over time.
@@ 6.6.4. Generalizability of downstream tasks
In our experiment, we chose personalized online advertisements as the target domain of downstream tasks due to their ubiquity, user familiarity, and sensitivity to modification of privacy data. Nevertheless, we expect that our approach can generalize to empowering users to experience the outcomes of their privacy behaviors in a wide range of downstream tasks, such as dynamic social media feeds, algorithmic decision-making, and news recommendations. For the next phase in the development of our Privacy Sandbox, we plan to add support for these additional downstream tasks, followed by the next rounds of deployment and experiments where we will assess our approach’s effectiveness in stimulating empathy, facilitating the acquisition of privacy knowledge, and promoting positive privacy behavior changes in these domains. In the end, upon study results validating its positive impacts, we will publicly release the Privacy Sandbox and promote its adoption through community outreach events for broader impacts.
@@ 7. Conclusion
Aiming to mitigate the gap of privacy paradox, we introduced an empathy-based approach that allows users to experience how privacy behaviors may alter system outcomes in a risk-free sandbox environment from the perspective of artificially generated personas. A user study with 15 participants confirmed the quality of the generated personas, validated the effectiveness of our approach in invoking user empathy and system outcome changes, and characterized the impact of users’ familiarity, personal experiences, and data consistency on their perceived realness and empathy toward these personas. Our findings offered design implications for implementing this approach in different downstream applications, with the aim of improving user privacy literacy and promoting behavior change.
@@ A.1. Prompts and few-shot learning examples for generating personas using our approach
{promptbox}
[title=Prompt to generate persona description] Return a realistic profile. This year is 2023. The income should be in dollars. The birthday should be in the MM/DD/YYYY format. The demographic of this person should represent the US population sample.
The generated profile should match the following guidance: ¡guidance¿.
Fit into the braces in the profile:
{First name} {Last name} is a {age ranging from 18 to 70 subject to continuous uniform distribution} {race} {gender} living in {real home address with street, city, state, and zip code}. {Pronoun} speaks {spoken language}. Pronoun’s education background is {educational background}. {Pronoun}’s date of birth is {date of birth}. {Pronoun} is a {occupation}, and the annual income is {income in dollar}. {marital status} {parental status} {detailed habits and preferences when using the computer, mobile phone, and the Internet}.
@@
[title=Few-shot learning example for the generation of persona description] Abigail Patel is a 32-year-old Asian American female living at 325 Main St, Newark, NJ 07102. She speaks English and her educational background includes a bachelor’s degree in Marketing. Abigail’s date of birth is 05/26/1991. She is currently working as a marketing manager, with an annual income of \$85,000. Abigail is married and has two children. She enjoys browsing social media and streaming movies on her mobile phone during her free time. When using her computer, she prefers using a wireless mouse and keyboard for easy navigation. On the internet, she likes to shop for clothes and read reviews before making a purchase.
{promptbox}
[title=Prompt to generate privacy attributes] ¡few-shot learning example¿
Given the profile: ¡persona¿.
Return the attributes in this format:
@@
{“first name”: “”, “last name”: “”, “age”: “”, “gender”: “”, “race”: “”, “street”: “”, “city”: “”, “state”: “”, ”zip code”: “”, “spoken language”: “”, “educational background”: “”, “birthday”: “”, “job”: “”, “income”: “”, “marital status”: “”, “parental status”: “”, “online behavior”: “”}
{promptbox}
@@
[title=Few-shot learning example for the generation of privacy attributes] Given the profile: Abigail Patel is a 32-year-old Asian American female living at 325 Main St, Newark, NJ 07102. She speaks English and her educational background includes a bachelor’s degree in Marketing. Abigail’s date of birth is 05/26/1991. She is currently working as a marketing manager, with an annual income of \$85,000. Abigail is married and has two children. She enjoys browsing social media and streaming movies on her mobile phone during her free time. When using her computer, she prefers using a wireless mouse and keyboard for easy navigation. On the internet, she likes to shop for clothes and read reviews before making a purchase.
Return the attributes in this format:
{“first name”: “Abigail”, “last name”: “Patel”, “age”: “32”, “gender”: “female”, “race”: “Asian American”, “street”: “325 Main St”, “city”: “Newark”, “state”: “NJ”, ”zip code”: “07102”, “spoken language”: “English”, “educational background”: “bachelor’s degree in Marketing”, “birthday”: “05/26/1991”, “job”: “marketing manager”, “income”: “85,000”, “marital status”: “married”, “parental status”: “has two children”, “online behavior”: “She enjoys browsing social media and streaming movies on her mobile phone during her free time. When using her computer, she prefers using a wireless mouse and keyboard for easy navigation. On the internet, she likes to shop for clothes and read reviews before making a purchase.”}
{promptbox}
@@
[title=Prompt for generating portrait image prompt] Given the profile description, output a descriptive prompt to generate a realistic human head portrait image, limit the word to 30 words: {persona description}
{promptbox}
@@
[title=Prompt to generate device and browser] Given the profile: {persona}, infer the browser and device the person uses:
{promptbox}
[title=Prompt to generate schedule] ¡few-shot learning example¿
@@
You are acting as a game event designer. Write daily events for this person: {persona description}. Show me a reasonable schedule for this person from {start_date} to {end_date}. The life in the period is similar to 2021. You can generate fake but reasonable data that is related to the profile. The start time of one day is 00:00:00. Generate events from 00:00:00 to 23:59:59 for each day.
Return a list of dict.
Output the following JSON format in plain text:
Never provide additional context.
{promptbox}
@@
[title=Few-shot learning example for the generation of schedule] “profile”: “Daniel Chan is a 30yearold Asian man living in Seattle, Washington with zip code 98101. He is a project manager, and the annual income is around one hundred and twenty thousand USD. He lives alone in a studio apartment and likes to keep his space clean and organized. In his free time, he enjoys playing video games and reading books on his Kindle. He also likes to use social media platforms such as Twitter and Reddit to keep up with the latest news and trends. ”
“location_history”: [ [“2023-06-05 00:00:00”, “2023-06-05 07:00:00”, “Home - 1420 5th Ave, Seattle, WA 98101”], [“2023-06-05 07:00:00”, “2023-06-05 08:30:00”, “Golds Gym - 1220 Howell St, Seattle, WA 98101”], [“2023-06-05 08:30:00”, “2023-06-05 09:00:00”, “Starbucks - 1125 4th Ave, Seattle, WA 98101”] ]
{promptbox}
[title=Prompt to generate browsing history] ¡few-shot learning example¿
Given the person’s profile: {persona description}, and the schedule: {schedule}, generate {number} browser history entries from {start_date} to {end_date}.
Never provide additional context.
{promptbox}
@@
[title=Few-shot learning example for the generation of browser history] “profile”: “John Smith is a 25yearold Caucasian male living at 123 Park Ave, New York, NY 10001. He speaks English and his educational background includes studying Computer Science and Data Analysis. John’s date of birth is 09/15/1998. He is currently a student, and his annual income is \$5000. John is single and does not have any children. He enjoys coding and exploring new technologies on his computer. On his mobile phone, he prefers using apps for productivity and staying up to date with the latest tech news. When using the Internet, he enjoys participating in online coding forums and watching tutorial videos to enhance his skills.”
@@
{promptbox}
[title=Prompt to generate social media post content] Provide ideas for this person to write posts (limit the word to 140 words) based on the profile and location history: {profile} {schedule}
The schedule is in the format of [[start time, end time, address]].
@@
{promptbox}
[title=Few-shot learning example for the generation of social media post content] “profile”: “Emily Rodriguez is a 46yearold Hispanic female living in 602 S Fairfax Ave, Los Angeles, CA 90036. She works as a nurse and earns an annual income of \$70,000. Emily is happily married with two children who are currently in college. In her free time, she enjoys reading and gardening. Emily prefers using her mobile phone for browsing social media and checking emails while using her laptop for work-related tasks. She is mindful of her online security and regularly updates her passwords and privacy settings.”
{promptbox}
[title=Prompt for the prompt to generate social media post image] Given the post {content}, output a descriptive prompt to generate a realistic life image, limit the word to 30 words:
A.2. Prompts for generating personas using baseline GPT
{promptbox}
[title=Prompt to generate persona description] Return a realistic profile. This year is 2023. The income should be in dollars. The birthday should be in the MM/DD/YYYY format. The demographic of this person should represent the US population sample.
The generated profile should match the following guidance: ¡guidance¿.
Fit into the braces in the profile:
{First name} {Last name} is a {age ranging from 18 to 70 subject to continuous uniform distribution} {race} {gender} living in {real home address with street, city, state, and zip code}. {Pronoun} speaks {spoken language}. Pronoun’s education background is {educational background}. {Pronoun}’s date of birth is {date of birth}. {Pronoun} is a {occupation}, and the annual income is {income in dollar}. {marital status} {parental status} {detailed habits and preferences when using the computer, mobile phone, and the Internet}.
Return the profile in only one paragraph.
{promptbox}
[title=Prompt to generate privacy attributes] Given the profile: ¡persona¿.
Return the attributes in this format:
{“first name”: “”, “last name”: “”, “age”: “”, “gender”: “”, “race”: “”, “street”: “”, “city”: “”, “state”: “”, ”zip code”: “”, “spoken language”: “”, “educational background”: “”, “birthday”: “”, “job”: “”, “income”: “”, “marital status”: “”, “parental status”: “”, “online behavior”: “”}
{promptbox}
[title=Prompt for generating portrait image] Generate a realistic human head portrait image
{promptbox}
[title=Prompt to generate device and browser] Generate the browser and device a person uses:
{promptbox}
[title=Prompt to generate schedule] You are acting as a game event designer. Write daily events for a persona. Show me a reasonable schedule for this person from {start_date} to {end_date}. The life in the period is similar to 2021. You can generate fake but reasonable data that is related to the profile. The start time of one day is 00:00:00. Generate events from 00:00:00 to 23:59:59 for each day.
Return a list of dict.
Output the following JSON format in plain text:
Never provide additional context.
{promptbox}
[title=Prompt to generate browsing history] Generate {number} browser history entries from {start_date} to {end_date}.
Never provide additional context.
{promptbox}
[title=Prompt to generate social media post content] Provide ideas for a person to write posts (limit the word to 140 words)
Never provide additional context.
{promptbox}
[title=Prompt to generate social media post image] Generate a realistic life image for social media posts
@@ Appendix B Participants’ demographic data
Persona ID Age Gender Education level Self Rated Digital Literacy
@@
Table 4. The demographic information of participants

### v2

@@ 6.6.3. Experiment design
The goal of the experiment design presented in this paper was to validate the feasibility of our approach in stimulating changes in system outcomes and invoking user empathy. As a result, the experiment did not aim to directly measure the extent to which users acquired privacy knowledge or the subsequent shifts in their privacy behavior as a consequence of our approach. Although the qualitative findings of the interviews with participants suggest a strong potential of our approach to achieve these two goals, more work is needed to validate such hypotheses in long term and in-the-wild context. Our study is an initial endeavor to enable users to empathize with generated personas, understand their privacy goals, and gain privacy insights by observing and correlating the system outcomes with privacy attributes. In the future, we plan to collect additional qualitative data when user evaluate the perceived realness across three groups of personas. This will help identify the key factors and issues that influence the perceived realness. We will also conduct pre-tests and post-tests with users between their usage of our system to assess their acquisition of privacy knowledge, as well as longitudinal deployment studies to measure changes in users’ privacy behaviors over time.
@@ 6.6.4. Generalizability of downstream tasks
In our experiment, we chose personalized online advertisements as the target domain of downstream tasks due to their ubiquity, user familiarity, and sensitivity to modification of privacy data. Nevertheless, we expect that our approach can generalize to empowering users to experience the outcomes of their privacy behaviors in a wide range of downstream tasks, such as dynamic social media feeds, algorithmic decision-making, and news recommendations. When acting as the generated personas with various personal profiles and social media posts, users can explore the association between privacy attributes and new friends or posts on social media. In addition, when browsing the recommended news presented to the generated personas with distinct browsing histories, online interests, and locations, users can also uncover the connections between these attributes with the recommended news. For the next phase in the development of our Privacy Sandbox, we plan to add support for these additional downstream tasks, followed by the next rounds of deployment and experiments where we will assess our approach’s effectiveness in stimulating empathy, facilitating the acquisition of privacy knowledge, and promoting positive privacy behavior changes in these domains. In the end, upon study results validating its positive impacts, we will publicly release the Privacy Sandbox and promote its adoption through community outreach events for broader impacts.
@@ 7. Conclusion
Aiming to mitigate the privacy gap between attitudes, goals, knowledge, and behavior, we introduced an empathy-based approach that allows users to experience how privacy attributes may alter system outcomes in a risk-free sandbox environment from the perspective of artificially generated personas. A user study with 15 participants confirmed the quality of the generated personas, validated the effectiveness of our approach in invoking user empathy and system outcome changes, and characterized the impact of users’ familiarity, personal experiences, and data consistency on their perceived realness and empathy toward these personas. Our findings offered design implications for implementing this approach in different downstream applications, with the aim of improving user privacy literacy.
Acknowledgements.
This work was supported in part by an AnalytiXIN Faculty Fellowship, an NVIDIA Academic Hardware Grant, a Google Cloud Research Credit Award, a Google Research Scholar Award, a Google PSS Privacy Research Award, NSF Grant 2341187, and NSF Grant 2326378. C. Chen and Y. Ye’s work was partially supported by the NSF under grants IIS-2321504, IS-2334193, IIS-2203262, IIS-2217239, CNS-2203261, and CMMI-2146076. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors. We would like to thank Tianshi Li and Dakuo Wang for useful discussions.
@@ A.1. Prompts and few-shot learning examples for generating personas using our approach
Prompt to generate persona description Return a realistic profile. This year is 2023. The income should be in dollars. The birthday should be in the MM/DD/YYYY format. The demographic of this person should represent the US population sample. The generated profile should match the following guidance: ¡guidance¿. Fit into the braces in the profile: {First name} {Last name} is a {age ranging from 18 to 70 subject to continuous uniform distribution} {race} {gender} living in {real home address with street, city, state, and zip code}. {Pronoun} speaks {spoken language}. Pronoun’s education background is {educational background}. {Pronoun}’s date of birth is {date of birth}. {Pronoun} is a {occupation}, and the annual income is {income in dollar}. {marital status} {parental status} {detailed habits and preferences when using the computer, mobile phone, and the Internet}.
@@
Few-shot learning example for the generation of persona description Abigail Patel is a 32-year-old Asian American female living at 325 Main St, Newark, NJ 07102. She speaks English and her educational background includes a bachelor’s degree in Marketing. Abigail’s date of birth is 05/26/1991. She is currently working as a marketing manager, with an annual income of \$85,000. Abigail is married and has two children. She enjoys browsing social media and streaming movies on her mobile phone during her free time. When using her computer, she prefers using a wireless mouse and keyboard for easy navigation. On the internet, she likes to shop for clothes and read reviews before making a purchase.
@@
Prompt to generate privacy attributes ¡few-shot learning example¿ Given the profile: ¡persona¿. Return the attributes in this format: {“first name”: “”, “last name”: “”, “age”: “”, “gender”: “”, “race”: “”, “street”: “”, “city”: “”, “state”: “”, ”zip code”: “”, “spoken language”: “”, “educational background”: “”, “birthday”: “”, “job”: “”, “income”: “”, “marital status”: “”, “parental status”: “”, “online behavior”: “”}
@@
Few-shot learning example for the generation of privacy attributes Given the profile: Abigail Patel is a 32-year-old Asian American female living at 325 Main St, Newark, NJ 07102. She speaks English and her educational background includes a bachelor’s degree in Marketing. Abigail’s date of birth is 05/26/1991. She is currently working as a marketing manager, with an annual income of \$85,000. Abigail is married and has two children. She enjoys browsing social media and streaming movies on her mobile phone during her free time. When using her computer, she prefers using a wireless mouse and keyboard for easy navigation. On the internet, she likes to shop for clothes and read reviews before making a purchase. Return the attributes in this format: {“first name”: “Abigail”, “last name”: “Patel”, “age”: “32”, “gender”: “female”, “race”: “Asian American”, “street”: “325 Main St”, “city”: “Newark”, “state”: “NJ”, ”zip code”: “07102”, “spoken language”: “English”, “educational background”: “bachelor’s degree in Marketing”, “birthday”: “05/26/1991”, “job”: “marketing manager”, “income”: “85,000”, “marital status”: “married”, “parental status”: “has two children”, “online behavior”: “She enjoys browsing social media and streaming movies on her mobile phone during her free time. When using her computer, she prefers using a wireless mouse and keyboard for easy navigation. On the internet, she likes to shop for clothes and read reviews before making a purchase.”}
@@
Prompt for generating portrait image prompt Given the profile description, output a descriptive prompt to generate a realistic human head portrait image, limit the word to 30 words: {persona description}
@@
Prompt to generate device and browser Given the profile: {persona}, infer the browser and device the person uses:
@@
Prompt to generate schedule ¡few-shot learning example¿ You are acting as a game event designer. Write daily events for this person: {persona description}. Show me a reasonable schedule for this person from {start_date} to {end_date}. The life in the period is similar to 2021. You can generate fake but reasonable data that is related to the profile. The start time of one day is 00:00:00. Generate events from 00:00:00 to 23:59:59 for each day. Return a list of dict. Output the following JSON format in plain text: { “start time”: ¡start moment of the event¿, “end time”: ¡start moment of the event¿, “event”: ¡event¿ } Never provide additional context.
@@
Few-shot learning example for the generation of schedule “profile”: “Daniel Chan is a 30yearold Asian man living in Seattle, Washington with zip code 98101. He is a project manager, and the annual income is around one hundred and twenty thousand USD. He lives alone in a studio apartment and likes to keep his space clean and organized. In his free time, he enjoys playing video games and reading books on his Kindle. He also likes to use social media platforms such as Twitter and Reddit to keep up with the latest news and trends. ” “location_history”: [ [“2023-06-05 00:00:00”, “2023-06-05 07:00:00”, “Home - 1420 5th Ave, Seattle, WA 98101”], [“2023-06-05 07:00:00”, “2023-06-05 08:30:00”, “Golds Gym - 1220 Howell St, Seattle, WA 98101”], [“2023-06-05 08:30:00”, “2023-06-05 09:00:00”, “Starbucks - 1125 4th Ave, Seattle, WA 98101”] ]
Prompt to generate browsing history ¡few-shot learning example¿ Given the person’s profile: {persona description}, and the schedule: {schedule}, generate {number} browser history entries from {start_date} to {end_date}. No browsing history between 00:00:00 and 07:00:00. The webpage title should reflect the content in the webpage url. The webpage be reasonable and related to the the schedule. Don’t add the address of the schedule to the webpage title. The datetime should be realistic and associated with the webpage content. The datetime second should not be 0. The datetime should be dispensed in one day. You can generate fake but reasonable data that is consistent with the profile and schedule. Output following list format in plain text:[[¡datetime¿, ¡webpage titile¿, ¡webpage url¿],] Never provide additional context.
@@
Few-shot learning example for the generation of browser history “profile”: “John Smith is a 25yearold Caucasian male living at 123 Park Ave, New York, NY 10001. He speaks English and his educational background includes studying Computer Science and Data Analysis. John’s date of birth is 09/15/1998. He is currently a student, and his annual income is \$5000. John is single and does not have any children. He enjoys coding and exploring new technologies on his computer. On his mobile phone, he prefers using apps for productivity and staying up to date with the latest tech news. When using the Internet, he enjoys participating in online coding forums and watching tutorial videos to enhance his skills.”
@@
Prompt to generate social media post content Provide ideas for this person to write posts (limit the word to 140 words) based on the profile and location history: {profile} {schedule} The schedule is in the format of [[start time, end time, address]]. Show me only num reasonable description in total between start_date and end_date to provide ideas. The life in the given time period is similar to 2021 so you can generate the description based on your current data. You should only return the list to me without any explanation message. You don’t need to use any real-time data, just generate reasonable and consistent data. You don’t need to generate descriptions that may be inappropriate, irrelevant, or offensive. You do not need to manipulate the data in a way that is specific to a given time period. The seconds in the time should not be 00, it should be the format like 15:23:12. Output the following JSON format in plain text: [{ ”time”: ¡time in string format¿, ”address”: ¡address where this person share the life¿, ”content”: ¡content¿, }] Never provide additional context.
Few-shot learning example for the generation of social media post content “profile”: “Emily Rodriguez is a 46yearold Hispanic female living in 602 S Fairfax Ave, Los Angeles, CA 90036. She works as a nurse and earns an annual income of \$70,000. Emily is happily married with two children who are currently in college. In her free time, she enjoys reading and gardening. Emily prefers using her mobile phone for browsing social media and checking emails while using her laptop for work-related tasks. She is mindful of her online security and regularly updates her passwords and privacy settings.”
Prompt for the prompt to generate social media post image Given the post {content}, output a descriptive prompt to generate a realistic life image, limit the word to 30 words:
A.2. Prompts for generating personas using baseline GPT
Prompt to generate persona description Return a realistic profile. This year is 2023. The income should be in dollars. The birthday should be in the MM/DD/YYYY format. The demographic of this person should represent the US population sample. The generated profile should match the following guidance: ¡guidance¿. Fit into the braces in the profile: {First name} {Last name} is a {age ranging from 18 to 70 subject to continuous uniform distribution} {race} {gender} living in {real home address with street, city, state, and zip code}. {Pronoun} speaks {spoken language}. Pronoun’s education background is {educational background}. {Pronoun}’s date of birth is {date of birth}. {Pronoun} is a {occupation}, and the annual income is {income in dollar}. {marital status} {parental status} {detailed habits and preferences when using the computer, mobile phone, and the Internet}.
Return the profile in only one paragraph.
Prompt to generate privacy attributes Given the profile: ¡persona¿. Return the attributes in this format: {“first name”: “”, “last name”: “”, “age”: “”, “gender”: “”, “race”: “”, “street”: “”, “city”: “”, “state”: “”, ”zip code”: “”, “spoken language”: “”, “educational background”: “”, “birthday”: “”, “job”: “”, “income”: “”, “marital status”: “”, “parental status”: “”, “online behavior”: “”}
Prompt for generating portrait image Generate a realistic human head portrait image
Prompt to generate device and browser Generate the browser and device a person uses:
Prompt to generate schedule
You are acting as a game event designer. Write daily events for a persona. Show me a reasonable schedule for this person from {start_date} to {end_date}. The life in the period is similar to 2021. You can generate fake but reasonable data that is related to the profile. The start time of one day is 00:00:00. Generate events from 00:00:00 to 23:59:59 for each day.
Return a list of dict.
Output the following JSON format in plain text:
Never provide additional context.
Prompt to generate browsing history
Generate {number} browser history entries from {start_date} to {end_date}.
Never provide additional context.
Prompt to generate social media post content
Provide ideas for a person to write posts (limit the word to 140 words)
@@ A.2. Prompts for generating personas using baseline GPT
Prompt to generate social media post image Generate a realistic life image for social media posts
@@ Appendix B Participants’ demographic data
Persona ID Age Gender Education level Digital Literacy
@@
Table 4. The demographic information of participants (The digital literacy is self rated)